OpenStudy (anonymous):

If the average life span of an elephant is 60 years, how many hours can you expect an elephant to live?

OpenStudy (anonymous):

There are 8765.81 hours in a year

OpenStudy (anonymous):

8765.81 * 60, and you have your answer.

OpenStudy (anonymous):

Your answer is 525948.6 hours

OpenStudy (tkhunny):

Multiply by 1. Works every time. \(60\;years\cdot\dfrac{365\;days}{1\;year}\cdot\dfrac{24\;hrs}{1\;day}\) -- Done.

OpenStudy (tkhunny):

You may wish to adjust for Leap Year, I suppose. You may also wish to adjust for the exact Solar Day. Blah, blah. Such adjustments won't amount to a hill of beans over 60 years.